Three force act on a partial and force it move with a constant velocity
V=(2m/s)i-(7m/s)j . Two of these force are F=(2N)i+(3N)j+(-2N)j and F2=(-5N)i+(8N)j+(-2N)k
Find the third force
1
Expert's answer
2018-04-24T09:48:09-0400
V=const => vector sum F1+F2+F3=0
=> sum i = 0, sum j = 0, sum k = 0
F3 = (xN)i + (yN)j + (zN)k
2N - 5N + xN = 0 => x = 3
3N - 2N + 8N +yN = 0 => y = -9
-2N + zN = 0 => z = 2
=> F3 = (3N)i + (-9N)j + (2N)k
Velocity vector does not affect